international space station makes 15.65 revolutions per day in its orbit around the earth. Assuming a circular orbit, how high is this satellite above the surface of the earth?
1
Expert's answer
2020-06-02T17:53:20-0400

The radius of ofbit is given by the expresson (see https://en.wikipedia.org/wiki/Geostationary_orbit#Derivation_of_geostationary_altitude):


"R = \sqrt[3]{\dfrac{\mu T^2}{4\pi^2}}"

where "\mu = 398.6\cdot 10^{3} km^3/s^2" is the the geocentric gravitational constant and "T = (15.65 rpd)^{-1} = (1.81\cdot 10^{-4} rps)^{-1} \approx5521s" is the period of revolution. Substituting values, obtain:


"R = \sqrt[3]{\dfrac{398.6\cdot 10^{3}\cdot 5521^2}{4\pi^2}} \approx 6752km"

Subtructing the Earth's equatorial radius, 6,378 kilometres, obtain the height of this satellite above the surface of the Earth:


"h = 6752-6378 = 374 km"

Answer. h = 374 km.